# Visual Description Grounding Reduces Hallucinations and Boosts Reasoning in LVLMs

### v2

@@ 5.2 Experimental Setup
Baselines. We compare VDGD against VCD, OPERA, Woodpecker, LRV, LURE, HALC (Chen et al., 2024a) and PAI (Liu et al., 2024d) hallucination mitigation techniques. Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla-greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7). We employ greedy decoding for all methods as we find no difference in performance on sampling. We generate image descriptions for VDGD using the VCD mitigation technique. For LVLMs, we employ LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, CogVLM, Qwen2-VL (Wang et al., 2024b) and CogVLM2 (Hong et al., 2024).
@@ 5.3 Results and Analysis
Quantitative results. Table 9 compares the performance of VDGD with other baselines. As we clearly see, VDGD not only outperforms all baselines but is also the only method that alleviates and improves performance on VaLLu. Results on other full benchmarks are provided in Appendix L.6, together with scores for other aspects beyond factuality, where we show that VDGD does not compromise the depth and expressivity of responses. Ablation results in Appendix C demonstrate that removing KL divergence leads to drop in performance. Additionally, using only the description prefix without the image leads to performance drops, confirming that VDGD enhances reasoning through grounding rather than making the model rely completely on the description alone.
Qualitative examples. Fig 10 compares responses from vanilla greedy decoding with VDGD. Grounding response generation results in more factual responses. Appendix B illustrates 2 failure cases of VDGD where incorrect descriptions can lead VDGD to hallucinate.
@@ 6 Conclusion, Limitations and Future Work
As part of future work, we would like to work on the current limitations of our work, which include (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
@@ 7 Reproducibility Statement
We provide our code here: https://anonymous.4open.science/r/VDGD-1E04/. All codes will be open-sourced upon paper acceptance, including all checkpoints. All experimental details, including training parameters and hyper-parameters, are provided in Section 5.3. The Appendix has more comprehensive details about our experimental setup.
@@ Appendix B VDGD Failure Cases
[width=]failure_cases.pdf
@@ Appendix C VDGD Ablation Results
To demonstrate the effectiveness of VDGD, we conduct ablation studies on its key components: (i) VDGD (- VCD): Image descriptions are generated without using Visual Contrastive Decoding (VCD), (ii) VDGD (- KLD): KLD-based sampling is removed, and only the description is appended, (iii) VDGD (- Image): VDGD is applied without the original image input; only the image description is appended to the prompt. Results are presented on the VDGD benchmark in Table 3. Removing KL Divergence leads to a substantial performance drop. We notice a similar phenomenon by removing the input image with the steepest drop. This shows that the input image is still essential for the model.
@@ Appendix D Correlation between GPT scores and actual metrics
Background of Experts in Human Evaluation. For the expert human evaluation of model generations, we engaged a group of 5 PhD students with at least four years of research experience in computer vision. These experts were tasked with evaluating the outputs of the models for accuracy, relevance, and factual correctness. To ensure thorough evaluation, they were provided with internet access to verify factual information when necessary. The evaluation process was conducted in an anonymized format to maintain objectivity, with the experts working independently and without prior knowledge of the model specifics. We selected 500 samples or the size of the benchmark, whichever was lower.
@@ Appendix I Model Details
From our in-depth analysis of the logit space, we see that language hallucinations in LVLMs have a unique characteristic – the probability of the most confident token in the vocabulary marginally decreased from the probability of the same token in the base LLM when prompted with the same prior context (the most confident token by the LVLM will also be the most confident token by the language-only base LLM as the token on a language hallucination). Thus, decoding-based mitigation methods like OPERA and VDD benefit from this unique property, and contrasting tokens help mitigate hallucinations.
@@
[width=]hallu_noisy_img_1.pdf
@@
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25 Vanilla-sampling 2.64 3.73 1.86 1.94 1.27 VDD 1.70 2.55 1.47 1.81 1.59 OPERA 2.60 3.75 2.04 1.89 1.30 Woodpecker 2.68 3.81 2.01 1.95 1.28 LRV 2.64 3.78 1.98 1.94 1.36 LURE 2.60 3.54 2.03 1.89 1.40 VDGD (ours) 2.38 3.39 2.16 2.46 1.34 LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25 Vanilla-sampling 2.57 3.65 2.01 1.92 1.22 VDD 1.94 2.99 1.55 1.65 1.05 OPERA 2.73 3.74 2.05 2.04 1.22 Woodpecker 2.74 3.87 2.05 2.03 1.29 LRV 2.76 3.82 2.10 2.03 1.28 LURE 2.72 3.76 2.03 2.05 1.27 VDGD (ours) 2.97 3.91 2.64 2.27 1.54 LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43 Vanilla-sampling 3.10 3.85 2.64 2.63 1.40 VDD 2.01 2.96 1.80 2.10 1.20 OPERA 3.10 4.01 2.62 2.60 1.42 Woodpecker 3.16 4.00 2.67 2.58 1.44 LRV 3.15 3.98 2.65 2.62 1.39 LURE 3.11 4.02 2.64 2.64 1.48 VDGD (ours) 3.51 4.04 3.16 2.87 1.58 mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28 Vanilla-sampling 2.72 3.60 2.18 1.91 1.24 VDD 2.16 2.95 1.80 1.85 1.26 OPERA 2.92 3.96 2.26 2.08 1.34 Woodpecker 2.89 3.90 2.23 2.02 1.24 LRV 2.90 3.94 2.19 2.05 1.28 LURE 2.92 3.89 2.24 2.06 1.32 VDGD (ours) 3.14 3.98 2.72 2.22 1.50 InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47 Vanilla-sampling 3.13 4.04 2.70 2.41 1.56 VDD 2.56 3.26 2.32 2.19 1.41 OPERA 3.14 4.20 2.65 2.32 1.46 Woodpecker 3.13 4.12 2.60 2.28 1.41 LRV 3.06 4.19 2.59 2.31 1.45 LURE 3.12 4.10 2.64 2.40 1.49 VDGD (ours) 3.57 4.37 3.45 2.65 1.68 CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55 Vanilla-sampling 3.26 4.21 2.83 2.40 1.55 VDD 2.80 3.62 2.63 2.10 1.58 OPERA 3.40 4.27 2.85 2.46 1.57 Woodpecker 3.44 4.30 2.91 2.45 1.59 LRV 3.42 4.26 2.88 2.42 1.54 LURE 3.39 4.20 2.78 2.41 1.52 VDGD (ours) 3.33 4.15 3.01 2.51 1.60 GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
@@
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00 Vanilla-sampling 1.82 2.82 1.38 1.56 1.06 VDD 1.70 2.67 1.22 1.40 0.95 OPERA 1.88 2.87 1.45 1.61 1.07 Woodpecker 1.89 2.88 1.48 1.60 1.06 LRV 1.83 2.85 1.40 1.57 1.08 LURE 1.91 2.90 1.52 1.63 1.09 VDGD (ours) 2.09 2.99 1.64 1.74 1.14 LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09 Vanilla-sampling 1.83 2.88 1.42 1.67 1.08 VDD 1.70 2.52 1.34 1.58 0.97 OPERA 2.18 3.08 1.62 1.75 1.22 Woodpecker 2.15 2.93 1.57 1.63 1.18 LRV 2.16 2.88 1.60 1.67 1.16 LURE 2.03 2.87 1.46 1.71 1.15 VDGD (ours) 2.30 3.17 1.85 1.86 1.20 LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97 Vanilla-sampling 1.80 2.63 1.50 1.78 0.96 VDD 1.68 2.57 1.36 1.68 0.94 OPERA 1.86 2.70 1.57 1.84 0.99 Woodpecker 1.90 2.74 1.63 1.89 1.04 LRV 1.88 2.72 1.61 1.85 1.02 LURE 1.92 2.73 1.65 1.86 1.03 VDGD (ours) 2.11 2.89 1.81 1.92 1.05 mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16 Vanilla-sampling 2.25 3.12 1.59 1.66 1.14 VDD 2.10 2.98 1.42 1.56 1.04 OPERA 2.35 3.23 1.67 1.74 1.18 Woodpecker 2.28 3.17 1.61 1.69 1.15 LRV 2.34 3.24 1.66 1.72 1.14 LURE 2.32 3.21 1.63 1.67 1.19 VDGD (ours) 2.48 3.36 1.82 1.78 1.23 InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63 Vanilla-sampling 2.97 3.76 2.17 2.45 1.64 VDD 2.76 3.54 2.03 2.26 1.57 OPERA 3.09 3.86 2.20 2.43 1.66 Woodpecker 3.07 3.83 2.19 2.46 1.68 LRV 3.11 3.88 2.24 2.49 1.70 LURE 3.10 3.89 2.22 2.48 1.71 VDGD (ours) 3.21 4.11 2.57 2.78 1.79 CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22 Vanilla-sampling 2.44 3.37 1.68 1.84 1.23 VDD 2.22 3.15 1.53 1.68 1.17 OPERA 2.56 3.47 1.76 1.88 1.27 Woodpecker 2.57 3.47 1.78 1.90 1.26 LRV 2.58 3.49 1.80 1.92 1.25 LURE 2.54 3.42 1.75 1.85 1.22 VDGD (ours) 2.65 3.56 1.92 1.97 1.29 GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
@@
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42 Vanilla-sampling 3.72 4.56 2.24 2.27 1.43 VDD 3.54 4.32 2.05 2.07 1.36 OPERA 3.80 4.63 2.29 2.35 1.50 Woodpecker 3.78 4.60 2.26 2.30 1.47 LRV 3.76 4.64 2.30 2.38 1.49 LURE 3.74 4.59 2.25 2.28 1.46 VDGD (ours) 3.92 4.78 2.52 2.42 1.54 LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32 Vanilla-sampling 3.68 4.39 2.38 2.04 1.36 VDD 3.38 4.09 2.35 1.87 1.10 OPERA 3.74 4.46 2.26 1.93 1.32 Woodpecker 3.46 4.53 2.34 1.97 1.38 LRV 3.53 4.42 2.39 1.96 1.27 LURE 3.33 4.41 2.37 1.91 1.29 VDGD (ours) 3.84 4.61 2.56 2.15 1.45 LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50 Vanilla-sampling 3.69 4.32 2.45 2.28 1.48 VDD 3.54 4.11 2.34 2.18 1.35 OPERA 3.80 4.47 2.59 2.46 1.57 Woodpecker 3.78 4.42 2.56 2.43 1.53 LRV 3.75 4.40 2.54 2.42 1.52 LURE 3.79 4.41 2.52 2.41 1.53 VDGD (ours) 3.90 4.53 2.79 2.51 1.58 mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33 Vanilla-sampling 3.80 4.45 2.51 2.12 1.34 VDD 3.62 4.27 2.35 1.98 1.24 OPERA 3.85 4.52 2.57 2.20 1.40 Woodpecker 3.84 4.56 2.60 2.24 1.42 LRV 3.82 4.54 2.59 2.22 1.41 LURE 3.86 4.58 2.61 2.25 1.45 VDGD (ours) 3.99 4.65 2.76 2.36 1.52 InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55 Vanilla-sampling 4.01 4.68 2.83 2.37 1.54 VDD 3.82 4.37 2.66 2.10 1.44 OPERA 4.13 4.76 2.92 2.42 1.59 Woodpecker 4.07 4.72 2.88 2.40 1.58 LRV 4.10 4.73 2.89 2.40 1.56 LURE 4.14 4.75 2.91 2.43 1.60 VDGD (ours) 4.28 4.89 3.12 2.54 1.66 CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44 Vanilla-sampling 3.73 4.38 2.74 2.61 1.42 VDD 3.58 4.27 2.64 2.50 1.37 OPERA 3.76 4.45 2.82 2.69 1.47 Woodpecker 3.75 4.44 2.79 2.64 1.45 LRV 3.80 4.51 2.86 2.70 1.49 LURE 3.82 4.53 2.88 2.73 1.52 VDGD (ours) 3.94 4.64 2.99 2.80 1.60 GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61

### v3

@@ 5.2 Experimental Setup
Baselines. We compare VDGD against VCD, OPERA, Woodpecker, LRV, LURE, HALC (Chen et al., 2024a) and PAI (Liu et al., 2024d) hallucination mitigation techniques. Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla-greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7). We employ greedy decoding for all methods as we find no difference in performance on sampling. For LVLMs, we employ LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, CogVLM, Qwen2-VL (Wang et al., 2024b) and CogVLM2 (Hong et al., 2024).
@@ 5.3 Results and Analysis
Quantitative results. Table 2 compares the performance of VDGD with other baselines. As we clearly see, VDGD not only outperforms all baselines but is also the only method that alleviates and improves performance on VaLLu. Results on other full benchmarks are provided in Appendix L.6, together with scores for other aspects beyond factuality, where we show that VDGD does not compromise the depth and expressivity of responses. Ablation results in Appendix C demonstrate that (i) removing KL divergence leads to drop in performance (ii) using only the description prefix without the image leads to performance drops confirming that VDGD enhances reasoning through grounding rather than making the model rely completely on the description alone, and (iii) more advanced captioning models can significantly enhance VDGD performance, while smaller captioning models provide competitive results with the added benefit of lower computational overhead.
Qualitative examples. Fig. 10 compares responses from vanilla greedy decoding with VDGD. Grounding response generation results in more factual responses. Appendix B illustrates 2 failure cases of VDGD where incorrect descriptions can lead VDGD to hallucinate.
@@ 6 Conclusion, Limitations and Future Work
As part of future work, we would like to work on (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix and (2) the requirement of prompting the LVLM twice.
@@ 7 Reproducibility Statement
We provide our code here: https://sreyan88.github.io/VDGD/. All experimental details, including training parameters and hyper-parameters, are provided in Section 5.2. The Appendix has more comprehensive details about our experimental setup.
@@ Appendix A Appendix
  12. 12.
    M: We provide details about computation cost for VDGD.
  13. 13.
    N: We provide results for VDGD with beam search.
  14. 14.
    O: We provide some further disccusion about related works.
@@ Appendix B VDGD Failure Cases
[width=]failure_cases_resized.pdf
@@ Appendix C VDGD Ablation Results
To demonstrate the effectiveness of VDGD, we conduct ablation studies on its key components: (i) VDGD (- VCD): Image descriptions are generated without using Visual Contrastive Decoding (VCD), (ii) VDGD (+ GPT-4 Captions): Image descriptions are generated using GPT-4-turbo-2024-04-09 instead of the LVLM itself. (iii) VDGD (+  Ramos et al. Captions): Image descriptions are generated using the small captioning model introduced by  Ramos et al. instead of the LVLM itself. (iv) VDGD (- KLD): KLD-based sampling is removed, and only the description is appended, (v) VDGD (- Image): VDGD is applied without the original image input; only the image description is appended to the prompt. Results are presented on the VDGD benchmark in Table 3. Removing KL Divergence leads to a substantial performance drop. We notice a similar phenomenon by removing the input image with the steepest drop. This shows that the input image is still essential for the model.
As we can clearly see, VDGD achieves a significant performance boost when using captions generated by GPT-4, highlighting the potential of higher-quality captions to enhance VDGD’s effectiveness. This is a case of implicit knowledge transfer from GPT-4 to VDGD through detailed descriptions. Additionally, VDGD delivers competitive results with a small-scale captioning model, illustrating an alternative, more computationally efficient approach to improving performance. These findings suggest that future advancements in both large and small captioning models could further enhance the performance and efficiency of VDGD.
@@ Appendix D Correlation between GPT scores and actual metrics
Background of Experts in Human Evaluation. For the expert human evaluation of model generations, we engaged a group of 5 PhD students with at least four years of research experience in computer vision. These experts were tasked with evaluating the outputs of the models for accuracy, relevance, and factual correctness. To ensure thorough evaluation, they were provided with internet access to verify factual information when necessary. The evaluation process was conducted in an anonymized format to maintain objectivity, with the experts working independently and without prior knowledge of the model specifics. We selected 500 samples or the size of the benchmark, whichever was lower. Scores presented in Table 4 are averaged across scores provided by the 5 students.
Table 4 presents scores in the format of original benchmark metrics / GPT Scores / Expert Human Scores. Remarkably, we show that our proposed GPT score has a higher correlation to Expert Human Score. Nonetheless, our GPT Score also has a high correlation with the original benchmark metrics.
[…]
0.9! Dataset Baseline LLaVA-1.5 mPLUG-Owl2 InternLM-X CogVLM2 Avg Correlation MME Vanilla 1501.45 / 3.54 / 3.91 1450.19 / 3.49 / 3.82 1712.00 / 3.75 / 4.15 1512.45 / 3.66 / 4.02 0.95 / 0.96 / 0.92 Woodpecker 1503.84/3.55/3.94 1508.87/3.63/3.98 1746.93 / 3.82 / 4.21 1789.63 / 3.92 / 4.33 VDGD 1698.23 / 3.70 / 4.12 1724.27 / 3.82 /4.24 1852.48 / 4.12 / 4.56 1848.87 / 4.09 / 4.34 MMMU-Val Vanilla 30.30 / 1.35 / 2.02 32.72 / 1.40 / 2.14 43.35 / 2.01 / 2.70 44.37 / 2.08 / 2.83 0.99 / 0.96 / 0.96 Woodpecker 33.12 / 1.44 / 1.83 36.35 / 1.61 / 1.79 47.58 / 2.12 / 2.60 48.98 / 2.14 / 2.53 VDGD 36.89 / 1.62 / 2.37 38.42 / 1.72 / 2.55 53.86 / 2.39 / 2.80 56.43 / 2.47 / 3.16 LLaVA-Bench Vanilla 65.42 / 3.10 / 3.49 69.37 / 3.17 / 3.60 72.21 / 4.08 / 4.22 73.76 / 4.17 / 4.19 0.96 / 0.97 / 0.95 Woodpecker 70.48 / 3.22 / 3.52 70.68 / 3.23 / 3.74 74.87 / 4.21 / 4.48 74.52 / 4.21 / 4.17 VDGD 71.03 / 3.47 / 4.04 70.57 / 3.38 / 3.91 77.43 / 4.45 / 4.72 78.03 / 4.50 / 4.73 MMVET Vanilla 31.15 / 2.50 /2.41 36.3 / 2.84 / 2.99 51.27 / 3.32 / 3.70 60.42 / 3.42 / 3.55 Woodpecker 33.42 / 2.64 / 2.88 47.26 / 3.06 / 3.29 51.67 / 3.40 / 3.81 65.78 / 3.50 / 3.75 0.97 / 0.98 / 0.91 VDGD 45.79 / 3.01 / 3.27 62.41 / 3.45 / 3.90 74.57 / 3.78 / 4.16 71.98 / 3.74 / 4.10
Table 4: Scores across 4 benchmarks in Original Metric / (Our proposed) GPT Score / (Our) Expert Human Score format. We also show the person correlation averaged across datasets. The correlation scores are in Original Metric-GPT Score / GPT Score-Expert Human Score / Original Metric-Expert Human Score format. GPT Scores show a higher correlation with Expert Human Scores.
@@ Appendix I Model Details
From our in-depth analysis of the logit space, we see that language hallucinations in LVLMs have a unique characteristic – the probability of the most confident token in the vocabulary marginally decreased from the probability of the same token in the base LLM when prompted with the same prior context (the most confident token by the LVLM will also be the most confident token by the language-only base LLM as the token on a language hallucination). Thus, decoding-based mitigation methods like OPERA and VCD benefit from this unique property, and contrasting tokens help mitigate hallucinations.
@@
[width=]hallu_noisy_img_1-compressed.pdf
@@
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25 Vanilla-sampling 2.64 3.73 1.86 1.94 1.27 VCD 1.70 2.55 1.47 1.81 1.59 OPERA 2.60 3.75 2.04 1.89 1.30 Woodpecker 2.68 3.81 2.01 1.95 1.28 LRV 2.64 3.78 1.98 1.94 1.36 LURE 2.60 3.54 2.03 1.89 1.40 VDGD (ours) 2.38 3.39 2.16 2.46 1.34 LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25 Vanilla-sampling 2.57 3.65 2.01 1.92 1.22 VCD 1.94 2.99 1.55 1.65 1.05 OPERA 2.73 3.74 2.05 2.04 1.22 Woodpecker 2.74 3.87 2.05 2.03 1.29 LRV 2.76 3.82 2.10 2.03 1.28 LURE 2.72 3.76 2.03 2.05 1.27 VDGD (ours) 2.97 3.91 2.64 2.27 1.54 LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43 Vanilla-sampling 3.10 3.85 2.64 2.63 1.40 VCD 2.01 2.96 1.80 2.10 1.20 OPERA 3.10 4.01 2.62 2.60 1.42 Woodpecker 3.16 4.00 2.67 2.58 1.44 LRV 3.15 3.98 2.65 2.62 1.39 LURE 3.11 4.02 2.64 2.64 1.48 VDGD (ours) 3.51 4.04 3.16 2.87 1.58 mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28 Vanilla-sampling 2.72 3.60 2.18 1.91 1.24 VCD 2.16 2.95 1.80 1.85 1.26 OPERA 2.92 3.96 2.26 2.08 1.34 Woodpecker 2.89 3.90 2.23 2.02 1.24 LRV 2.90 3.94 2.19 2.05 1.28 LURE 2.92 3.89 2.24 2.06 1.32 VDGD (ours) 3.14 3.98 2.72 2.22 1.50 InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47 Vanilla-sampling 3.13 4.04 2.70 2.41 1.56 VCD 2.56 3.26 2.32 2.19 1.41 OPERA 3.14 4.20 2.65 2.32 1.46 Woodpecker 3.13 4.12 2.60 2.28 1.41 LRV 3.06 4.19 2.59 2.31 1.45 LURE 3.12 4.10 2.64 2.40 1.49 VDGD (ours) 3.57 4.37 3.45 2.65 1.68 CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55 Vanilla-sampling 3.26 4.21 2.83 2.40 1.55 VCD 2.80 3.62 2.63 2.10 1.58 OPERA 3.40 4.27 2.85 2.46 1.57 Woodpecker 3.44 4.30 2.91 2.45 1.59 LRV 3.42 4.26 2.88 2.42 1.54 LURE 3.39 4.20 2.78 2.41 1.52 VDGD (ours) 3.33 4.15 3.01 2.51 1.60 GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
@@
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00 Vanilla-sampling 1.82 2.82 1.38 1.56 1.06 VCD 1.70 2.67 1.22 1.40 0.95 OPERA 1.88 2.87 1.45 1.61 1.07 Woodpecker 1.89 2.88 1.48 1.60 1.06 LRV 1.83 2.85 1.40 1.57 1.08 LURE 1.91 2.90 1.52 1.63 1.09 VDGD (ours) 2.09 2.99 1.64 1.74 1.14 LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09 Vanilla-sampling 1.83 2.88 1.42 1.67 1.08 VCD 1.70 2.52 1.34 1.58 0.97 OPERA 2.18 3.08 1.62 1.75 1.22 Woodpecker 2.15 2.93 1.57 1.63 1.18 LRV 2.16 2.88 1.60 1.67 1.16 LURE 2.03 2.87 1.46 1.71 1.15 VDGD (ours) 2.30 3.17 1.85 1.86 1.20 LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97 Vanilla-sampling 1.80 2.63 1.50 1.78 0.96 VCD 1.68 2.57 1.36 1.68 0.94 OPERA 1.86 2.70 1.57 1.84 0.99 Woodpecker 1.90 2.74 1.63 1.89 1.04 LRV 1.88 2.72 1.61 1.85 1.02 LURE 1.92 2.73 1.65 1.86 1.03 VDGD (ours) 2.11 2.89 1.81 1.92 1.05 mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16 Vanilla-sampling 2.25 3.12 1.59 1.66 1.14 VCD 2.10 2.98 1.42 1.56 1.04 OPERA 2.35 3.23 1.67 1.74 1.18 Woodpecker 2.28 3.17 1.61 1.69 1.15 LRV 2.34 3.24 1.66 1.72 1.14 LURE 2.32 3.21 1.63 1.67 1.19 VDGD (ours) 2.48 3.36 1.82 1.78 1.23 InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63 Vanilla-sampling 2.97 3.76 2.17 2.45 1.64 VCD 2.76 3.54 2.03 2.26 1.57 OPERA 3.09 3.86 2.20 2.43 1.66 Woodpecker 3.07 3.83 2.19 2.46 1.68 LRV 3.11 3.88 2.24 2.49 1.70 LURE 3.10 3.89 2.22 2.48 1.71 VDGD (ours) 3.21 4.11 2.57 2.78 1.79 CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22 Vanilla-sampling 2.44 3.37 1.68 1.84 1.23 VCD 2.22 3.15 1.53 1.68 1.17 OPERA 2.56 3.47 1.76 1.88 1.27 Woodpecker 2.57 3.47 1.78 1.90 1.26 LRV 2.58 3.49 1.80 1.92 1.25 LURE 2.54 3.42 1.75 1.85 1.22 VDGD (ours) 2.65 3.56 1.92 1.97 1.29 GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
@@
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42 Vanilla-sampling 3.72 4.56 2.24 2.27 1.43 VCD 3.54 4.32 2.05 2.07 1.36 OPERA 3.80 4.63 2.29 2.35 1.50 Woodpecker 3.78 4.60 2.26 2.30 1.47 LRV 3.76 4.64 2.30 2.38 1.49 LURE 3.74 4.59 2.25 2.28 1.46 VDGD (ours) 3.92 4.78 2.52 2.42 1.54 LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32 Vanilla-sampling 3.68 4.39 2.38 2.04 1.36 VCD 3.38 4.09 2.35 1.87 1.10 OPERA 3.74 4.46 2.26 1.93 1.32 Woodpecker 3.46 4.53 2.34 1.97 1.38 LRV 3.53 4.42 2.39 1.96 1.27 LURE 3.33 4.41 2.37 1.91 1.29 VDGD (ours) 3.84 4.61 2.56 2.15 1.45 LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50 Vanilla-sampling 3.69 4.32 2.45 2.28 1.48 VCD 3.54 4.11 2.34 2.18 1.35 OPERA 3.80 4.47 2.59 2.46 1.57 Woodpecker 3.78 4.42 2.56 2.43 1.53 LRV 3.75 4.40 2.54 2.42 1.52 LURE 3.79 4.41 2.52 2.41 1.53 VDGD (ours) 3.90 4.53 2.79 2.51 1.58 mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33 Vanilla-sampling 3.80 4.45 2.51 2.12 1.34 VCD 3.62 4.27 2.35 1.98 1.24 OPERA 3.85 4.52 2.57 2.20 1.40 Woodpecker 3.84 4.56 2.60 2.24 1.42 LRV 3.82 4.54 2.59 2.22 1.41 LURE 3.86 4.58 2.61 2.25 1.45 VDGD (ours) 3.99 4.65 2.76 2.36 1.52 InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55 Vanilla-sampling 4.01 4.68 2.83 2.37 1.54 VCD 3.82 4.37 2.66 2.10 1.44 OPERA 4.13 4.76 2.92 2.42 1.59 Woodpecker 4.07 4.72 2.88 2.40 1.58 LRV 4.10 4.73 2.89 2.40 1.56 LURE 4.14 4.75 2.91 2.43 1.60 VDGD (ours) 4.28 4.89 3.12 2.54 1.66 CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44 Vanilla-sampling 3.73 4.38 2.74 2.61 1.42 VCD 3.58 4.27 2.64 2.50 1.37 OPERA 3.76 4.45 2.82 2.69 1.47 Woodpecker 3.75 4.44 2.79 2.64 1.45 LRV 3.80 4.51 2.86 2.70 1.49 LURE 3.82 4.53 2.88 2.73 1.52 VDGD (ours) 3.94 4.64 2.99 2.80 1.60 GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61